A chloro compound A on reduction with Zn-Cu and alcohol gives the hydrocarbon with five carbon
atoms. When A is dissolved in ether and treated with sodium 2,2,5,5-tetramethylhexane(B) is formed.
Write the structures of A and B.
1
Expert's answer
2013-07-09T08:22:04-0400
Subjecting the haloalkanes (R-X) to a reducing medium such as Zn/Cu couple + ethanol will give a direct replacement of the halogen (X)
with an hydrogen atom:
R-X + 2[H] = R-H + HX
If alkyl halide react with sodium metal, the alkyl halides attach themselves to each other at the place where the halogen was (the Wurtz
reaction). Products of the Wurtz reaction are always symmetric:
2R-X + 2Na = R-R + 2NaX
Thus, the compound B ( 2,2,5,5-tetramethylhexane ) is aresult of the Wurtz reaction. This compound is symmetric, so one half of
it is R:

B:
CH3-C(CH3)2-CH2-CH2-C(CH3)2-CH3

R:
CH3-C(CH3)2-CH2-

Thus, the compound A is corresponding chloro-derivative of R:

CH3-C(CH3)2-CH2-Cl (1-Chloro-2,2-dimethylpropane, Neopentyl chloride).